OpenStudy (anonymous):

Find the limit as x is approaching 0 of cot(3x)sin(6x) Please tell me where I am going wrong in my work, which I will post below in the comments.

OpenStudy (anonymous):

So this is the question: \[\lim_{x \rightarrow 0} \cot(3x)\sin(6x)\] *I made it a fraction, preparing it for L'Hoptals. \[\sin(6x)/\tan(3x)\] *I do L'Hopitals, taking derivative! \[6\cos(6x)/3\csc^2(3x)\] *simplify \[2\cos(6x)/\csc^2(3x)\] No I dunno what in the world to do next :(

OpenStudy (anonymous):

What in the world... WolframAlpha took the derivative of sin(6x)/tan(3x), and somehow got: \[2\cos(3x)^2\cos(6x)\] ??? How wolfram alpha!

myininaya (myininaya):

o \[\lim_{x \rightarrow 0}\frac{\cos(3x)}{\sin(3x)}\sin(6x) \frac{\frac{1}{6x}}{\frac{1}{3x}} \cdot \frac{\frac{6}{1}}{\frac{3}{1}}\] \[\lim_{x \rightarrow 0}\cos(3x) \cdot \frac{1}{\frac{\sin(3x)}{3x}} \cdot \frac{\sin(6x)}{6x} \cdot \frac{6}{3}\] \[\cos(3 \cdot 0) \cdot \frac{1}{1} \cdot 1 \cdot \frac{6}{3}\] \[1 \cdot 1 \cdot 1 \cdot 2=2\]

OpenStudy (anonymous):

Okay? so i am still confused? I don't know why you did what you did. From the first step...

myininaya (myininaya):

so you don't know this: \[\lim_{u \rightarrow 0}\frac{\sin(u)}{u}=1 \] ?

OpenStudy (anonymous):

No. What is that? i've never seen that beforee!

myininaya (myininaya):

thats weird they teach that before l'hospital's rule

myininaya (myininaya):

like way before we can prove it by squeeze thm

OpenStudy (anonymous):

Yeah i know squeeze theroem. But i need to know how to do it the normal, intedned way. L'hoptials.

myininaya (myininaya):

why is that the normal way?

myininaya (myininaya):

the way i showed you is easiest

myininaya (myininaya):

i just multiplied by 1/1 so i could apply that lim(u->0) sin(u)/u=1

OpenStudy (anonymous):

Wait, I also didn't remember that cos/sin = cot. About this sinu/u thing? Are there any others?

OpenStudy (anonymous):

I thought cot=1/tan...

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\cos(x)-1}{x}=0\]

myininaya (myininaya):

we can prove that multiplying top and bottom by the conjugate of the top and then applying lim(x->0) sin(x)/x=1

OpenStudy (anonymous):

I need to do this using lhoptials. We are learning how to use lhoptals. The whole point of what I'm doing is to learn the application of the rule.

myininaya (myininaya):

\[\cot(x)=\frac{adj}{opp} ; \frac{\cos(x)}{\sin(x)}=\frac{\frac{adj}{hyp}}{\frac{opp}{hyp}}=\frac{adj}{opp}\] \[=>\cot(x)=\frac{\cos(x)}{\sin(x)}\]

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin(6x)}{\tan(3x)}\] so we have 0/0 we can apply l'hospital rule \[\lim_{x \rightarrow 0}\frac{6 \cos(6x)}{3 \sec^2(3x)}\] now just plug in 0

myininaya (myininaya):

\[\frac{6(1)}{3(1^2)}=2\]

OpenStudy (anonymous):

Okay so i get were the cos/sin comes from. SO let me try it this way now: \[\cos(3x)/\sin(3x)*sin(6x)\] Simplify \[\sin(3x)\cos(3x)\] L'hopitals \[[\cos(3x)(3)][-\sin(3x)(3)]\] am i on the right track?

myininaya (myininaya):

why are you making it so hard i used l'hospital's rule with just one round of it and then i was able to use direct substitution

OpenStudy (anonymous):

K, looking at yours, sec.

myininaya (myininaya):

wow wolfram does have a freaking long answer lol

myininaya (myininaya):

so unnecessary though

OpenStudy (anonymous):

Hey, look at my first step! I did exactly what you did! SO i've been diogn it right?

OpenStudy (anonymous):

I just simplified the 6/3 = 2/1

myininaya (myininaya):

you do know \[\frac{d}{dx}(\tan(x))=\sec^2(x) \text{ and \not } \csc^2{x} \text{ iright ?}\]

OpenStudy (anonymous):

Wait, so I know what a cos curve looks like. So i can deduce cos(0) np. So the numerator is 6. But I don't know what sec^2(0) is.. I mean I know it's 1 because I saw the answer. But I don't know the sec^2 curve by heart. Is the only solution to just memorize it like I have for cos tan sin?

myininaya (myininaya):

\[\sec(x)=\frac{1}{\cos(x)}\]

myininaya (myininaya):

if cos(x)=1, then sec(x) is also 1

OpenStudy (anonymous):

I see... So cot = 1/tan AND cos/sin? Didn't realize it had to.. Okay well this makes sense now, I can't thank you enough!

myininaya (myininaya):

there are so many trig identities

OpenStudy (anonymous):

Sorry to ask but can i get help with another? Creating the question now..